What volume of 3.10 M HCL can be made from 100.0 g of HCL
1
Expert's answer
2020-04-19T07:08:37-0400

n(HCl)=m(HCl)/Mr(HCl)=100/35.5=2.82 mol;

C(HCl)=n(HCl)/V(HCl);

V(HCl)=n(HCl)/C(HCl)=m(HCl)/Mr(HCl)*C(HCl)=100/(35.5*3.10)=0.91 L;

Answer: V(HCl)= 0.91 L